OpenStudy (anonymous):

solve over the interval [0,360) 5cot^2 theta -cot theta = 2

OpenStudy (anonymous):

Try to put : \[x=\cot\theta\] And find the solution to the new equation of x :)

OpenStudy (anonymous):

i got (1±√41)/10 but i dont know where to go from there

OpenStudy (anonymous):

its 1 plus or minus the square root of 41 over 10

OpenStudy (anonymous):

OK ! After that you have : \[\cot\theta=\frac{1+\sqrt{41}}{10}\text{ or }\cot\theta=\frac{1-\sqrt{41}}{10}.\] Solve each equation and you will get the solution .

OpenStudy (anonymous):

so there is no need for me to convert it to tan before i solve to find the solution?

OpenStudy (anonymous):

The answer would be 0.74, because -0.5 will not fit in the interval.